OpenStudy (anonymous):

how to solve -1=6-n . like : -1=6-n -1-6=-n -7=-n then ??

OpenStudy (mathmate):

-1=6-n add 1 to each side 1-1=1+6-n => 0=7-n add n to each side n+0=7-n+n n=7

OpenStudy (mathmate):

Or from -7=-n multiply by -1 on each side 7=n which is the same as n=7